------------------------------------------------------------ Date: Mon Aug 14 00:22:15 2000, Shift: 16:00-24:00 Supervisor: JHL, with: B3 Log book#4 pages 81 - 83 Shift Summary: ------------------------------------------------------------ 1. Beam - Stored and cogged for 5.5 hours for this shift. - One store runs about 4 hours with intensities of 5.5 *10^9 for Blue and 2.8*10^9 for Yellow. 2. Rates - Event Rate: 13 Events/sec at the beginning of the store and 9 Events/sec at the near the store - Trigger rate: Trigger 4: 10.3/sec Trigger 6: 1.7/sec 3. Vertex - BB and ZDC vertices are Centered ~ 0. - sigma of BB vertex: 43cm (beginning of store) - 54 cm (end of store) 4. Running conditions: - MRS at 90, FS at 5 degree - All Global detectors, MTPC1,MTPC2,T1, Hodo's... - TPC's seem to be happy. Especially T1 shows clean multiple tracks - D1, D2 are on. D5 (4KG A-pol) is on since run 2238 . 5. Data - run 2235: 24323 collisions - run 2236: 26143 collisions - run 2237: 15763 collisions - run 2238: still running (expected to have more than 30k collisions) 6. Problems - All HV MF are okay. - T1 Tripped once. - D5 was "turned on" in the previous shift, But Dana found out that no current was reading in P.S. When magnet is turned on, one has to make sure that the output value should read same as set point value in the monitor. ------------------------------------------------------------
This archive was generated by hypermail 2b29 : Mon Aug 14 2000 - 00:27:35 EDT